OpenStudy (anonymous):

Given that y = (x –5)(2x+5)^0.5 , find dy/dx

OpenStudy (anonymous):

Use product then chain rule: \[\Large \begin{array}{l} y = (x - 5){(2x + 5)^{1/2}}\\ u'v + uv'\\ u = (x - 5)\\ u' = 1\\ v = {(2x + 5)^{1/2}}\\ v' = a'[b] \cdot b'\\ a = {b^{1/2}}\\ a' = \frac{1}{2}{b^{ - 1/2}}\\ b = 2x + 5\\ b' = 2\\ v' = \frac{1}{2}{(2x + 5)^{ - 1/2}} \cdot 2 = {(2x + 5)^{ - 1/2}}\\ y' = 1 \cdot {(2x + 5)^{1/2}} + (x - 5) \cdot {(2x + 5)^{ - 1/2}} \end{array}\] And simplify.